# appendixReferences

### v1

@@ 5 Conclusion and Discussion
Limitations and Discussion. The noise in data is a challenge for the models based on conditional generation. In the experiment, we find the noise of object attributes in grids will influence the selection accuracy of generative solvers like RAISE and Transformer on 2$\times$2Grid. The candidate sets can provide clearer supervision in training to reduce the impact of noise. Deep latent variable models (DLVMs) can potentially handle noise in RPMs since RAISE works well in Center and O-IC with noisy attributes like Rotation. In future works, exploring appropriate ways to reduce the influence of noise is the key to realizing generative abstract reasoning in more complicated scenes. However, for generative solvers that do not rely on candidate sets or are completely unsupervised, the noise level of datasets is a problem worth discussing. Too much noise will make a problem have a large number of solutions (e.g., PGM Barrett et al. (2018)), such data may not be proper for validating the generative reasoning ability of models. In Appendices B.2 and D.1, we conduct an initial experiment and discussion on the impact of noise, but a more systematic and in-depth study will be carried out in the follow-up works.
@@ A.1 Reformulation of the posterior distribution
The posterior distribution $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ can be computed from the prior $q(r^{c}|\bm{z}_{S}^{c})$ and the likelihood $p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})$ in Bayes’ theorem:
@@
According to Equation 4, $p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})$ is a Gaussian distribution with the mean $h\left(\bm{Z}^{c};\psi_{r^{c}}\right)$ and the standard deviation $\sigma_{z}$. Therefore, Equation 13 can be expressed as
@@
where $\text{D}(\bm{z}_{T}^{c})$ is the size of $\bm{z}_{T}^{c}$. In practice, neural networks output unnormalized logits $\bm{\tilde{l}}_{1:K}^{c}$ instead of the probabilities $\bm{\tilde{\pi}}_{1:K}^{c}$. Therefore, we use the logarithmic version of Equation 14:
@@
$C(\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ is not related to $r^{c}$ and thus will not influence the results of the softmax operation on $r^{c}$. RAISE ignores $C(\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ in Equation 15 and predicts the unnormalized logits via
@@ A.2 Derivation of the ELBO
Considering the generative process and inference processes
@@
$q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})$ $=\prod_{c=1}^{C}q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})\prod_{s\in S}q_{\theta}(\bm{z}_{s}^{c}|\bm{x}_{s})\prod_{t\in T}q_{\theta}(\bm{z}_{t}^{c}|\bm{x}_{t}),$
@@
$=\sum_{t\in T}\mathbb{E}_{q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})}\big{[}\log p_{\varphi}(\bm{x}_{t}|\bm{z}_{t})\big{]}-\sum_{c=1}^{C}\mathbb{E}_{q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})}\left[\log\frac{\prod_{t\in T}q_{\theta}(\bm{z}_{t}^{c}|\bm{x}_{t})}{p_{\psi}(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})}\right]$
@@
Since the encoder is shared between the generative and inference processes, we have $q_{\theta}(\bm{z}_{s}^{c}|\bm{x}_{s})=p_{\theta}(\bm{z}_{s}^{c}|\bm{x}_{s})$ and then
@@
$\mathcal{L}$ $=\underbrace{\sum_{t\in T}\mathbb{E}_{q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})}\big{[}\log p_{\varphi}(\bm{x}_{t}|\bm{z}_{t})\big{]}}_{\text{$\mathcal{L}_{\text{rec}}$}}-\underbrace{\sum_{c=1}^{C}\mathbb{E}_{q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})}\left[\log\frac{\prod_{t\in T}q_{\theta}(\bm{z}_{t}^{c}|\bm{x}_{t})}{p_{\psi}(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})}\right]}_{\text{$\mathcal{R}_{\text{pred}}$}}$ (22)
@@ A.3 Monte Carlo Estimator of the ELBO
To compute $\mathcal{L}$ for a given RPM $(\bm{x}_{S},\bm{x}_{T})$, we sample the latent variables $\bm{\tilde{r}}$, $\bm{\tilde{z}}_{S}$, and $\bm{\tilde{z}}_{T}$ from the variatonal posterior $q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})$ through
@@ A.3.2 Concept Regularizer
$\mathcal{R}_{\text{pred}}$ $=\sum_{c=1}^{C}\mathbb{E}_{q_{\theta}(\bm{z}_{T}^{c}|\bm{x}_{T})}\left[\mathbb{E}_{q_{\theta}(\bm{z}_{S}^{c}|\bm{x}_{S})}\left[\mathbb{E}_{q_{\phi,\psi}(r^{c}|\bm{\tilde{z}}_{S}^{c},\bm{\tilde{z}}_{T}^{c})}\left[\log\frac{\prod_{t\in T}q_{\theta}(\bm{z}_{t}^{c}|\bm{x}_{t})}{p_{\psi}(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})}\right]\right]\right]$ (25)
@@
Based on the supervision of $r^{c}$, RAISE can quickly approach the real distribution $p_{\star}(r^{c}|\bm{x})$ after the early learning stage. That is, the predicted rule distribution is close to the real one provided as the supervision, which is related to the image matrix itself rather than conditional on the latent concepts. Therefore, we assume that there is $q_{\phi,\psi}(r^{c}|\bm{\tilde{z}}_{S}^{c},\bm{\tilde{z}}_{T}^{c})\approx p_{\star}(r^{c}|\bm{x})$ after a few learning epochs, by which we can move the inner expectation on $r^{c}$ to the front. In this way, the inner expectation becomes the KL divergence between Gaussian distributions with a closed-form solution, which reduces the noise in the sampling process.
@@ A.3.3 Rule Regularizer
$\mathcal{R}_{\text{rule}}$ $=\sum_{c=1}^{C}\mathbb{E}_{q_{\theta}(\bm{z}_{T}^{c}|\bm{x}_{T})}\left[\mathbb{E}_{q_{\theta}(\bm{z}_{S}^{c}|\bm{x}_{S})}\left[\mathbb{E}_{q_{\phi,\psi}(r^{c}|\bm{\tilde{z}}_{S}^{c},\bm{\tilde{z}}_{T}^{c})}\left[\log\frac{q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})}{p_{\phi}(r^{c}|\bm{z}_{S}^{c})}\right]\right]\right]$ (26)
@@ A.3.4 ELBO
Ignoring the constants $C_{\text{rec}}$ and $C_{\text{pred}}$, the approximation of ELBO is
@@ Appendix B Datasets
Figure 5: Samples from the seven image configurations of RAVEN. In (a)-(g), the top panels are 3$\times$3 RPMs where the target images are highlighted in red boxes; the middle panels are candidate sets; and the bottom panels show attribute-specific rules in the RPMs.
@@ B.1 RAVEN and I-RAVEN
  1. 1.
    Constant: an attribute stays unchanged in rows;
  2. 2.
    Progress: an attribute increases or decreases equidistantly in rows;
  3. 3.
    Arithmetic: the attribute of the third image is equal to the sum or difference of that in the first two images;
  4. 4.
    Distribution Three: the attribute in rows are three fixed values in different orders.
Each configuration contains 6000 training samples, 2000 validation samples, and 2000 test samples. RAVEN provides eight candidate images and attribute-level annotations of rules for RPMs.
[…]
@@ B.2 Attribute Noise of RAVEN and I-RAVEN
Figure 6: The illustration of attribute noise. (a) is an RPM from 2$\times$2Grid; (b) is the candidate set; (c) and (d) display two possible types of noise in the RPM. In this case, the image is correct as long as there are two pentagons of the correct size. The color, rotation, and position of objects will not influence the correctness of the image.
RAVEN and I-RAVEN introduce noise to some attributes to increase the complexity of problems. In Center, L-R, U-D, and O-IC, the rotation of objects is the noise attribute. That is, we can keep the rotation of objects constant in rows or randomly sample the rotations. Object grids in O-IG, 2$\times$2Grid, and 3$\times$3Grid have more complex noise. Figure 6 displays the noise of object grids, including the noise of object attributes (i.e., objects in Figure 6c can have different colors and rotations), and the noise of object positions (Figure 6d). The candidate set ensures that only one candidate image is the correct answer, which can provide clear supervision of the rule. To explore the influence of noise, we remove the noise of object attributes from O-IG, 2$\times$2Grid, and 3$\times$3Grid, keep the noise of object positions, and generate three novel configurations O-IG-Uni, 2$\times$2Grid-Uni, and 3$\times$3Grid-Uni.
@@ C.1 RAISE
    where $K$ is the number of atomic rules to select from. Since the row and column representations are concatenated as the input, the input size of the network is 128.
  • $h(\bm{Z}^{c};\psi_{k})$. The network is a fully convolutional network, which predicts the means of target latent concepts from the representation matrix $\bm{Z}^{c}$:
@@
    $h$ adopts convolutional layers with 3$\times$3 kernels, stride 1, and padding 1 to keep the shape of the 3$\times$3 representation matrix. Here the global knowledge set $\psi_{1:K}$ stores $K$ groups of parameters for $h$, each of which represents an atomic rule.
@@ C.3 Baselines
Table 3: The learning rate of ANP on each configuration of RAVEN/I-RAVEN.
@@
Table 4: Hyperparameters of CLAP. We give the number of concepts (nc), weights in the ELBO ($\beta_{t}$, $\beta_{f}$, and $\beta_{TC}$), and concept standard deviation $\sigma_{z}$ of each configuration.
@@ D.1 Bottom-Right Answer Selection
Table 5: The accuracy (%) of selecting bottom-right answers on O-IG-Uni, 2$\times$2Grid-Uni, and 3$\times$3Grid-Uni. They are generated by removing the noise in object attributes.
@@
We generate three configurations by removing the noise in object attributes to analyze the influence of noise attributes. The results in Table 5 indicate that RAISE achieves the highest accuracy on the new configurations. Intuitively, the number of correct answers to an RPM increases when complex noise attributes exist. In this case, the provided candidate set contains one correct answer and seven distractors, which can act as clear supervision in model training. However, since RAISE and Transformer are trained without the aid of candidate sets, it is challenging to catch rules from noisy RPMs with multiple potential answers. Therefore, RAISE and Transformer have a more significant accuracy improvement. Overall, the experimental results show that reducing noise can bring significant improvements for the models trained without distractors in candidate sets (such as Transformer and RAISE). Moreover, RAISE only requires 20% rule annotations in this experiment to accomplish atomic rule learning via low-noise samples.
@@ D.2 Answer Selection at Arbitrary Position
Figure 7: Selection accuracy at arbitrary positions on I-RAVEN. Each plot contains the selection accuracy of RAISE (purple), Transformer (orange), CLAP (green), ANP (blue), and LGPP (black) at arbitrary positions. The x-axis indicates the number of candidates, and the y-axis is the selection accuracy.
In this subsection, we give additional experimental results of answer generation at arbitrary positions. Figure 8 provides the detailed results of arbitrary-position answer generation for all seven configurations of RAVEN, for example, the prediction results when $|T|=1$ (Figure 8a) and $|T|=2$ (Figure 8b). In the visualization results, RAISE can generate high-quality predictions when $|T|=1$ and $|T|=2$. The performance of Transformer varies significantly among different configurations. Transformer predicts accurate answers on Center, while the predictions on 3$\times$3Grid deviate a lot from the ground truth images. In most cases, ANP, LGPP, and CLAP tend to generate incorrect images. Figure 7 provides the selection accuracy on I-RAVEN with different numbers of targets ($|T|=1,2$) and different sizes of candidate sets ($N_{c}=1,3,7,15$) as additional results. And we can conduct further analysis through the selection accuracy with test errors in Tables 7 and 8, where RAISE clearly outperforms other baseline models on all image configurations of RAVEN and I-RAVEN.

### v3

@@ 5 Conclusion and Discussion
Limitations and Discussion. The noise in data is a challenge for the models based on conditional generation. In the experiment, we find that the noise of object attributes in grids will influence the selection accuracy of generative solvers like RAISE and Transformer on 2$\times$2Grid. The candidate sets can provide clearer supervision in training to reduce the impact of noise. Deep latent variable models (DLVMs) can potentially handle noise in RPMs since RAISE works well on Center and O-IC with noisy attributes like Rotation. In future works, exploring appropriate ways to reduce the influence of noise is the key to realizing generative abstract reasoning in more complicated scenes. For generative solvers that do not rely on candidate sets or are completely unsupervised, whether using datasets with large amounts of noise benefits the acquisition of generative abstract reasoning ability is worth exploring since the noise can make a generative problem have numerous solutions (e.g., PGM Barrett et al. (2018)). In Appendices B.2 and D.1, we conduct an initial experiment and discussion on the impact of noise, but a more systematic and in-depth study will be carried out in the follow-up works. Some recent neural approaches attempt to solve similar systematic generalization problems Rahaman et al. (2021); Lake & Baroni (2023). We provide a discussion on the Bayesian and neural approaches of concept learning in Appendix E.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No.62176060) and the Program for Professor of Special Appointment (Eastern Scholar) at Shanghai Institutions of Higher Learning.
@@ A.1 Reformulation of the posterior distribution
According to Bayes’ theorem, the posterior distribution of rule indicators $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ is the product of the conditional prior $q(r^{c}|\bm{z}_{S}^{c})$ and the likelihood $p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})$:
@@
Considering that $p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})$ is an isotropic Gaussian $\mathcal{N}\left(h\left(\bm{Z}^{c};\psi_{r^{c}}\right),\sigma_{z}^{2}\bm{I}\right)$, Equation 13 becomes
@@
where $\text{D}(\bm{z}_{T}^{c})$ is the size of $\bm{z}_{T}^{c}$. In practice, RAISE predicts unnormalized logits $\bm{\tilde{l}}_{1:K}^{c}$ instead of the probabilities $\bm{\tilde{\pi}}_{1:K}^{c}$. Therefore, we use the logarithmic version of Equation 14:
@@
Since the constant $C(\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ in Equation 15 will not influence the results of normalization, RAISE ignores the constant term and predicts the unnormalized logits via
@@ A.2 Derivation of the ELBO
Considering the generative and inference processes
@@
$q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})$ $=\prod_{c=1}^{C}\left(q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})\prod_{s\in S}q_{\theta}(\bm{z}_{s}^{c}|\bm{x}_{s})\prod_{t\in T}q_{\theta}(\bm{z}_{t}^{c}|\bm{x}_{t})\right),$
@@
$=\sum_{t\in T}\mathbb{E}_{q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})}\big{[}\log p_{\varphi}(\bm{x}_{t}|\bm{z}_{t})\big{]}-\sum_{c=1}^{C}\mathbb{E}_{q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})}\left[\log\frac{q_{\theta}(\bm{z}_{T}^{c}|\bm{x}_{T})}{p_{\psi}(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})}\right]$
@@
Since the encoder is shared between the generative and inference processes, we have $q_{\theta}(\bm{z}_{s}^{c}|\bm{x}_{s})=p_{\theta}(\bm{z}_{s}^{c}|\bm{x}_{s})$ and
@@
$\mathcal{L}$ $=\underbrace{\sum_{t\in T}\mathbb{E}_{q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})}\big{[}\log p_{\varphi}(\bm{x}_{t}|\bm{z}_{t})\big{]}}_{\text{$\mathcal{L}_{\text{rec}}$}}-\underbrace{\sum_{c=1}^{C}\mathbb{E}_{q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})}\left[\log\frac{q_{\theta}(\bm{z}_{T}^{c}|\bm{x}_{T})}{p_{\psi}(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})}\right]}_{\text{$\mathcal{R}_{\text{pred}}$}}$ (22)
@@ A.3 Monte Carlo Estimator of the ELBO
For a given RPM problem $(\bm{x}_{S},\bm{x}_{T})$, we sample the latent variables $\bm{\tilde{r}}$, $\bm{\tilde{z}}_{S}$, and $\bm{\tilde{z}}_{T}$ from the variatonal posterior $q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})$ to compute the ELBO:
@@ A.3.2 Concept Regularizer
$\mathcal{R}_{\text{pred}}$ $=\sum_{c=1}^{C}\mathbb{E}_{q_{\theta}(\bm{z}_{T}^{c}|\bm{x}_{T})}\left[\mathbb{E}_{q_{\theta}(\bm{z}_{S}^{c}|\bm{x}_{S})}\left[\mathbb{E}_{q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})}\left[\log\frac{q_{\theta}(\bm{z}_{T}^{c}|\bm{x}_{T})}{p_{\psi}(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})}\right]\right]\right]$ (25)
@@
Trained with rule annotations, RAISE can quickly approach the real distribution $p_{\star}(r^{c}|\bm{x})$ provided in rule annotations after the early learning stage. Therefore, we regard the real distribution as the predicted rule distribution, which is related to the matrix rather than conditional on the latent concepts. That is, we assume that samples from $p_{\star}(r^{c}|\bm{x})$ are similarly distributed to those from $q_{\phi,\psi}(r^{c}|\bm{\tilde{z}}_{S}^{c},\bm{\tilde{z}}_{T}^{c})$ after a few learning epochs. By replacing the $q_{\phi,\psi}(r^{c}|\bm{\tilde{z}}_{S}^{c},\bm{\tilde{z}}_{T}^{c})$ to $p_{\star}(r^{c}|\bm{x})$, we move the inner expectation on $r^{c}$ to the front. In this way, the inner expectation becomes the KL divergence between Gaussians and has a closed-form solution, reducing the additional noise in the sampling process.
@@ A.3.3 Rule Regularizer
$\mathcal{R}_{\text{rule}}$ $=\sum_{c=1}^{C}\mathbb{E}_{q_{\theta}(\bm{z}_{T}^{c}|\bm{x}_{T})}\left[\mathbb{E}_{q_{\theta}(\bm{z}_{S}^{c}|\bm{x}_{S})}\left[\mathbb{E}_{q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})}\left[\log\frac{q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})}{p_{\phi}(r^{c}|\bm{z}_{S}^{c})}\right]\right]\right]$ (26)
@@ A.3.4 ELBO
Ignoring the constant terms $C_{\text{rec}}$ and $C_{\text{pred}}$, the approximation of ELBO is
@@ Appendix B Datasets
Figure 5: Different configurations of RAVEN. In each figure, the top panel is an RPM where the target images are highlighted in red boxes; the middle panel is a candidate set with eight candidate images; and the bottom panel shows the attribute-changing rules in the RPM.
@@ B.2 Attribute Noise of RAVEN and I-RAVEN
Figure 6: The illustration of attribute noise. (a) is an RPM from 2$\times$2Grid; (b) is the candidate set; (c) and (d) visualize two possible types of noise in the RPM. In this case, the image is correct as long as there are two pentagons of the correct size. The color, rotation, and position of objects will not influence the correctness of the image.
RAVEN and I-RAVEN introduce noise to some attributes to increase the complexity of problems. In Center, L-R, U-D, and O-IC, the rotation of objects is the noise attribute. We can keep objects unchanged in rows or make random rotations. Figure 6 displays the noise of object grids on O-IG, 2$\times$2Grid, and 3$\times$3Grid, including the noise of object attributes (i.e., objects in Figure 6c can have different colors and rotations), and the noise of object positions (Figure 6d). The candidate set ensures that only one candidate image is the correct answer. To explore the influence of noise on selection accuracy, we remove the noise of object attributes from object grids, keep the noise of object positions, and generate three configurations O-IG-Uni, 2$\times$2Grid-Uni, and 3$\times$3Grid-Uni.
@@ C.1 RAISE
    where $K$ is the number of atomic rules. Since the row and column representations are concatenated as the input, the input size of the network is 128.
  • $h(\bm{Z}^{c};\psi_{k})$. This network is a fully convolutional network, which predicts the means of target latent concepts from the representation matrix $\bm{Z}^{c}$:
@@
    $h$ adopts convolutional layers with 3$\times$3 kernels, stride 1, and padding 1 to keep the shape of the 3$\times$3 representation matrix. The global knowledge set $\psi_{1:K}$ stores $K$ learnable parameters of $h$, which represents $K$ atomic rule respectively.
@@ C.3 Baselines
Table 3: Learning rates of ANP on RAVEN/I-RAVEN.
@@
Table 4: Hyperparameters of CLAP. We give the number of concepts, weights in the ELBO ($\beta_{t}$, $\beta_{f}$, and $\beta_{TC}$), and standard deviation $\sigma_{z}$ on RAVEN/I-RAVEN.
@@ D.1 Bottom-Right Answer Selection
Table 5: The accuracy (%) of selecting bottom-right answers on O-IG-Uni, 2$\times$2Grid-Uni, and 3$\times$3Grid-Uni.
@@
Table 6: The accuracy (%) of selecting bottom-right answers on different configurations (i.e., Center, L-R, etc) of RAVEN/I-RAVEN. In this table, RAISE is trained without the supervision of rule annotations (-aux) to illustrate the abstract reasoning ability in the unsupervised training setting. The table displays the average results of ten trials.
Models Average Center L-R U-D O-IC O-IG 2$\times$2Grid 3$\times$3Grid
RAISE (-aux) 54.5/67.7 30.2/56.6 47.9/80.8 87.0/94.9 96.9/99.2 56.9/83.9 30.4/30.5 32.0/27.8
We generate new configurations by removing the noise in object attributes to analyze the influence of noise attributes. As shown in Table 5, RAISE achieves the highest accuracy on all three configurations. When we introduce more noise to RPMs, the number of solutions that follow the correct rules will increase. In this case, the provided candidate set with one correct answer and seven distractors can act as clear supervision in model training. Without the assistance of candidate sets in training, it is challenging to catch rules from noisy RPMs with multiple potential solutions. Therefore, RAISE and Transformer have significant accuracy improvements on configurations with less noise attributes. Overall, the experimental results show that reducing noise can bring significant improvements for the models trained without distractors in candidate sets (such as Transformer and RAISE). RAISE only requires 20% rule annotations to learn atomic rules from low-noise samples.
We also provide the selection accuracy of unsupervised RAISE in Table 6. The average accuracy of unsupervised RAISE lies between the unsupervised arbitrary-generation baselines (i.e., LGPP, ANP, CLAP, and Transformer) and the powerful generative RPM solvers trained with full rule annotations (i.e., GCA, ALANS, and PrAE).
@@ D.2 Answer Selection at Arbitrary Position
Figure 7: Selection accuracy at arbitrary positions on I-RAVEN. Each plot contains the selection accuracy of RAISE (purple), Transformer (orange), CLAP (green), ANP (blue), and LGPP (black). The x-axis is the number of candidates, and the y-axis is the selection accuracy.
In this section, we give additional results for arbitrary-position answer generation. Figure 8 provides the detailed results of arbitrary-position answer generation for all seven configurations of RAVEN, for example, the prediction results when $|T|=1$ (Figure 8a) and $|T|=2$ (Figure 8b). In the visualization results, RAISE can generate high-quality predictions when $|T|=1$ and $|T|=2$. The performance of Transformer varies significantly among different configurations. Transformer predicts accurate answers on Center, while the predictions on 3$\times$3Grid deviate significantly from the ground truth images. In most cases, ANP, LGPP, and CLAP tend to generate incorrect images. Figure 7 provides the selection accuracy on I-RAVEN with different numbers of target images ($|T|=1,2$) and different numbers of distractors in candidate sets ($N_{c}=1,3,7,15$). We can make further analysis through the selection accuracy with test errors in Tables 8 and 9, where RAISE outperforms other baseline models on all image configurations of RAVEN and I-RAVEN.